OpenStudy (anonymous):

5^(1+x) = 2^(1-x) Solve using logss

OpenStudy (anonymous):

start with \[(1+x)\ln(5)=(1-x)\ln(2)\] and then algebra to solve for \(x\)

OpenStudy (anonymous):

very similar problem just posted here http://openstudy.com/study#/updates/5098605de4b02ec0829c7ebe

OpenStudy (anonymous):

@satellite73 thank you!

OpenStudy (anonymous):

yw